# Atomic Calibration of LLMs in Long-Form Generations

### v1

@@ Abstract
Large language models (LLMs) often suffer from hallucinations, posing significant challenges for real-world applications. Confidence calibration, which estimates the underlying uncertainty of model predictions, is essential to enhance the LLMs’ trustworthiness. Existing research on LLM calibration has primarily focused on short-form tasks, providing a single confidence score at the response level (macro calibration). However, this approach is insufficient for long-form generations, where responses often contain more complex statements and may include both accurate and inaccurate information. Therefore, we introduce atomic calibration, a novel approach that evaluates factuality calibration at a fine-grained level by breaking down long responses into atomic claims. We classify confidence elicitation methods into discriminative and generative types and demonstrate that their combination can enhance calibration. Our extensive experiments on various LLMs and datasets show that atomic calibration is well-suited for long-form generation and can also improve macro calibration results. Additionally, atomic calibration reveals insightful patterns in LLM confidence throughout the generation process.
@@ 1 Introduction
While large language models (LLMs), such as Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023), and GPT models (OpenAI, 2022), excel in various tasks, they still struggle with faithfulness and reliability issues. LLMs often suffer from hallucinations, generating factually inaccurate content and misleading responses (Zhang et al., 2023b; Huang et al., 2023), which limits their application in high-risk real-world scenarios (Hu et al., 2023). To address this, confidence calibration aims to estimate the underlying uncertainty of model predictions and reflect the true likelihood of correctness (Guo et al., 2017). A calibrated model is crucial for real-world applications, as it allows us to determine the extent to which we can trust models’ predictions (Zhu et al., 2023; Mahaut et al., 2024).
@@
Most existing work on LLM calibration focuses on short-form QA tasks (Jiang et al., 2021; Tian et al., 2023; Zhu et al., 2023; Ulmer et al., 2024). These studies often use datasets like TriviaQA (Joshi et al., 2017) and Natural Questions (Joshi et al., 2017), where answers typically have fewer than 10 words. However, in real-world applications, responses to user queries are often much longer (Zhang et al., 2024), sometimes extending to hundreds or even thousands of words. In such cases, the quality of LLM responses is not simply binary (correct or incorrect), as answers may include both accurate and inaccurate statements. Previous work on long-form calibration has primarily focused on response-level calibration (Zhang et al., 2024; Huang et al., 2024) (which we term macro calibration). This approach provides an overall confidence estimation for the entire response (as shown in the upper part of Figure 1). However, it is insufficient for long-form generation, as it cannot adequately capture the model’s fine-grained uncertainty across multiple factual statements.
In this work, we introduce the concept of atomic calibration, which evaluates calibration at the fine-grained level of atomic claims (as illustrated in the lower part of Figure 1). We focus on the aspect of factuality, since hallucination are a widely recognized issue of LLMs (Zhang et al., 2023b; Huang et al., 2023) and the factuality of statements can be objectively determined. Accurate factual calibration is crucial for mitigating the hallucination problem (Mahaut et al., 2024). To perform atomic calibration, we decompose long responses into self-contained atomic claims and evaluate the calibration for all atomic claims. Various confidence elicitation methods are studied to provide accurate confidence estimation. We provide a detailed investigation into long-form generation calibration, covering 7 LLMs and 3 datasets. The main contributions of our work are as follows:
  • We propose the novel concept of atomic calibration. Unlike macro calibration, which provides a single response-level confidence score, atomic calibration offers a more fine-grained analysis of model calibration, making it better suited for long-form generation (§2).
  • We introduce a categorization of confidence elicitation methods for long-form generation into two types: discriminative and generative, estimating the model’s intrinsic and external confidence, respectively. We examine five confidence elicitation methods and assess their effectiveness for atomic calibration (§3).
  • We show that atomic calibration is particularly well-suited for long-form generation, and that macro calibration can be enhanced by incorporating atomic calibration results. We also find the two types of confidence elicitation methods are complementary and their fusion can provide better calibration results. Atomic calibration also enables deeper analysis, revealing patterns in confidence and calibration changes throughout the generation process (§5).
@@ 2 Atomic Calibration
For a language model $\mathcal{M}$, let $x\in\mathcal{X}$ represent the response generated by model $\mathcal{M}$ for a query $q$, and let $y\in\mathcal{Y}_{t}$ denote the corresponding label, where $\mathcal{Y}_{t}\subseteq[0,1]$ indicates a quality score for a specific task $t\in T$. Note that unlike multiple-choice and short-form questions, which typically only focus on the correctness of the answer, the tasks in $T$ involve various aspects, such as factuality, coherence, creativity, etc.
We define a probability prediction function $f:\mathcal{X}\to\Delta^{\lvert\mathcal{Y}_{t}\rvert}$, where $\Delta^{\lvert\mathcal{Y}_{t}\rvert}$ represents the $\lvert\mathcal{Y}_{t}\rvert$-dimensional simplex. In this context, $f(x)_{y}$ denotes the probability assigned to the label $y$ given the generated output $x$. In the remaining of this work, we focus specifically on calibrating for factuality. In this case, $\mathcal{Y}$ represent $\mathcal{Y}_{t}$ when $t$ corresponds to factuality. $\mathcal{Y}$ takes values in [0, 1], indicating the factuality level of a response. Based on the calibration concepts discussed in (Guo et al., 2017), we define the calibration of each response as follows:
@@
$\mathbb{P}(y\mid f(x)_{y}=\beta)=\beta,\quad\forall\beta\in\Delta^{\lvert\mathcal{Y}\rvert}.$
@@
$\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta_{i},\quad\forall\beta_{i}\in\Delta^{\lvert\mathcal{Y}_{i}\rvert}.$
@@ 3 Confidence Elicitation Methods
To calculate the confidence scores for model outputs in long-form generation, we define two types of confidence elicitation methods: generative and discriminative. Generative methods assume that the consistency between different generation samples provides a reliable estimation of model uncertainty. The more frequently certain facts are covered by the samples, the higher the model’s confidence in those facts. In contrast, discriminative methods assess uncertainties by asking the model itself. This is motivated by the findings that models tend to perform better on discriminative tasks (Saunders et al., 2022), and thus they may already possess the capability to estimate the confidence of their own outputs in a discriminative manner.
We first generate one response $x$ as the answer to the query $q$. Then, $x$ is broken into atomic claims $C$. Following (Min et al., 2023; Wei et al., 2024; Zhao et al., 2024), each atomic claim contains a single piece of information and must be self-contained. For generative methods, we sample an additional set of responses $K$, and compare them against the original response $x$. For each atomic claim in $C$, we assign it a confidence score.
@@ 3.1 Generative Methods
Gen-Binary. The basic assumption here is that if a fact is frequently conveyed when sampled multiple times, the model is considered “confident” about that fact. For an atomic claim $c_{i}$ in $C$, we utilize a natural language inference model $\mathcal{M}_{\mathrm{NLI}}$ to examine whether $c_{i}$ is supported or not supported by each of the additional samples. Let $K_{s}$ be the set of samples supporting $c_{i}$. Then, the confidence in $c_{i}$ is calculated as:
$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K\rvert}$
Gen-Multi. Gen-Multi assumes that the model is more confident in facts that are consistently expressed. Unlike Gen-Binary, it further divides the “not supported” class $K_{ns}$ into “conflict” ($K_{c}$) if the fact is presented differently in the sample, and “not mentioned” ($K_{nm}$) if the fact is not mentioned in the sample. We then calculate the confidence by only considering supporting and conflicting samples:
$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K_{s}\rvert+\lvert K_{c}\rvert}$
@@ 3.2 Discriminative Methods
Dis-Single. Following (Kadavath et al., 2022; Tian et al., 2023), we directly ask the model whether one single atomic claim is true or false. The probability the model assigns to token “True” ($P(true)$) in its generation is viewed as the confidence. As each atomic claim is judged individually, one advantage of this method is that there is no cross-claim influences when the model makes confidence judgments.
@@
Dis-Rating. Instead of using $P(true)$, in Dis-Rating, we directly prompt the model to assign a numerical value representing its confidence in the atomic claim $c_{i}$. A score of 0 indicates no confidence, while 10 represents maximum confidence. An alternative approach is to use semantic expressions ranging from “Very Uncertain” to “Very Confident”. However, Tian et al. (2023) show that LLMs express uncertainty as effectively, or even more effectively, using numerical values rather than words.
@@ 3.3 Confidence Fusion Strategies
Since the generative and discriminative methods provide different perspectives of confidence estimation, it is natural to explore fusion methods that combine different confidence estimates for better calibration. After exploration, we consider four simple but effective options: (1) MinConf selects the minimum of the two confidence values, providing a conservative estimate; (2) HMean calculates the harmonic mean, providing a balanced fusion that weights lower confidence values more heavily, to help prevent overconfident predictions; (3) ProdConf multiplies the confidences, ensuring a high final confidence only when both estimates are strong, thereby effectively penalizing discrepancies between the two estimates; and (4) WAvg computes the weighted average, with a hyper-parameter weight, prioritizing the more reliable estimate and allowing flexibility to adjust the influence of each confidence source based on prior knowledge or validation results.
4 Evaluation Metrics
4.1 Atomic Calibration
In atomic calibration, we utilize FActScore (Min et al., 2023) to measure the atomic claims’ factuality. For each claim, we have its binary factuality label and a continuous confidence score. Since this corresponds to standard calibration scenarios, we follow the conventions from previous work (Kuhn et al., 2022; Zhu et al., 2023; Tian et al., 2023) and use Expected Calibration Error (ECE) (Naeini et al., 2015), Brier Score (Brier, 1950), and AUROC to assess calibration. Details of the three metrics are in Appendix A.
4.2 Macro Calibration
In macro calibration, factuality scores for responses are represented as continuous values from 0 to 1, therefore, the atomic calibration metrics mentioned above are not directly applicable. Apart from the Spearman Correlation, we propose two novel metrics UCCE and QCCE that extend ECE to continuous factuality scores.
Spearman Correlation. Following (Zhang et al., 2024; Huang et al., 2024), we calculate Spearman Correlation to assess whether samples with higher factuality have correspondingly higher confidence scores. Compared to Pearson Correlation, it focuses on assessing the rank correlation, is robust to outliers and does not require that data is in normal distribution.
Uniform Continuous Calibration Error (UCCE). UCCE evaluates how well a model’s continuous predictions align with the true values by dividing the predicted values into bins of equal size. This metric assesses the model’s calibration across the range of predictions. The calibration error is computed by comparing the mean predicted and true values within each bin.
Let $\hat{y}_{i}$ represent the predicted value for the $i$-th sample, $y_{i}$ the true value for the $i$-th sample, $B_{m}$ the set of samples falling into the $m$-th bin, $M$ the number of bins, and $N$ the total number of samples. The UCCE is computed as follows:
$\text{UCCE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\frac{\sum_{i\in B_{m}}\hat{y}_{i}}{|B_{m}|}-\frac{\sum_{i\in B_{m}}y_{i}}{|B_{m}|}\right|.$
UCCE is particularly effective for scenarios where predicted values are uniformly distributed.
Quantile Continuous Calibration Error (QCCE). QCCE uses a similar approach to UCCE but utilizes quantile binning, ensuring that each bin contains an equal number of samples. Consequently, each bin has a size of $\lvert B_{m}\rvert=\frac{N}{M}$. QCCE is advantageous for skewed or non-uniform distributions of predicted values, as it avoids sparsity in any bin by ensuring a balanced representation across all bins. The QCCE is calculated as follows:
QCCE $=\frac{1}{M}\sum_{m=1}^{M}\left|\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{y}_{i}-\frac{1}{|B_{m}|}\sum_{i\in B_{m}}y_{i}\right|$
$=\frac{1}{N}\sum_{m=1}^{M}\left|\sum_{i\in B_{m}}\hat{y}_{i}-\sum_{i\in B_{m}}y_{i}\right|.$
@@ 5.1 Experiment Setup
Models. We utilize seven LLMs from three model families with varying sizes: Llama3 Instruct (8B and 70B) (Meta, 2024), Mistral Instruct (7B and 8x7B) (Jiang et al., 2023), and Qwen2 Instruct (7B, 52B-A14B, and 72B) (Yang et al., 2024).
@@
Atomic Facts Generation and Verification. For all three datasets, we apply a FActScore-based (Min et al., 2023) factuality assessment approach. We first use GPT-4o to decompose the entire response into atomic facts, with each fact containing only a single piece of information. These atomic facts are then verified using GPT-4o, cross-referenced with evidence from Wikipedia and Google Search. The detailed prompts for generating atomic facts are provided in Appendix B.
Confidence Elicitation. We use P(true) (Kadavath et al., 2022), Self-Rating (Tian et al., 2023), and Semantic Entropy (SE) (Kuhn et al., 2022) as the baseline confidence elicitation methods. They are all calculated in response-level. For Gen-Binary, we apply the Llama3-8B-Instruct for better NLI performance. In WAvg, we assign a weighting 0.7 to generative methods and 0.3 to discriminative ones, based on validation results.
@@
SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
@@
P(true) 30.2 45.1 46.0 18.9 16.3 16.8 40.5 25.7 26.1
Self-Rating 40.5 38.7 39.3 21.5 14.1 14.3 50.2 18.6 19.0
SE 42.1 37.4 38.0 23.0 13.5 13.8 52.0 17.8 18.2
@@
Dis-Rating 73.8 25.7 25.7 34.1 2.9 2.8 71.7 3.6 3.8
Dis-Single 58.0 27.2 27.2 20.9 8.7 8.4 55.9 14.4 14.1
Gen-Binary 79.8 5.6 5.6 52.7 3.0 3.0 70.0 7.8 7.2
Gen-Multi 71.4 35.8 36.0 37.5 11.6 11.8 62.6 22.0 24.1
@@
P(true) 32.8 44.5 44.9 22.0 16.7 17.0 41.2 24.3 25.2
Self-Rating 42.3 37.1 37.5 26.1 14.5 14.7 52.0 18.1 19.3
SE 44.1 36.5 36.7 28.0 13.8 14.1 53.4 17.4 18.5
@@
Dis-Rating 55.0 41.4 41.5 40.8 4.4 4.2 60.4 16.9 16.7
Dis-Single 70.3 16.0 16.0 32.8 8.8 9.0 65.3 10.4 9.7
Gen-Binary 74.9 8.5 8.0 64.1 2.5 2.5 73.9 10.3 10.3
Gen-Multi 60.7 38.7 38.7 49.6 11.9 12.6 65.6 26.2 27.5
@@
P(true) 33.5 45.0 44.8 28.3 11.2 11.6 35.4 12.7 12.9
Self-Rating 48.2 24.3 23.9 36.7 6.9 7.0 48.0 9.8 9.5
SE 49.8 22.9 22.6 38.9 6.5 6.4 49.2 8.9 9.2
@@
Dis-Rating 63.0 40.7 41.0 29.9 4.4 4.4 54.0 9.1 8.9
Dis-Single 52.8 19.8 19.3 30.9 4.9 5.1 60.4 5.3 5.6
Gen-Binary 72.4 5.4 5.7 67.6 2.0 1.8 72.2 6.5 6.4
Gen-Multi 43.1 38.8 38.3 52.1 11.4 11.8 63.2 21.6 23.9
Table 2: Response Calibration Results. All the numbers are in percentages.

### v2

@@ Abstract
Large language models (LLMs) often suffer from hallucinations, posing significant challenges for real-world applications. Confidence calibration, as an effective indicator of hallucination, is thus essential to enhance the trustworthiness of LLMs. Prior work mainly focuses on short-form tasks using a single response-level score (macro calibration), which is insufficient for long-form outputs that may contain both accurate and inaccurate claims. In this work, we systematically study atomic calibration, which evaluates factuality calibration at a fine-grained level by decomposing long responses into atomic claims. We further categorize existing confidence elicitation methods into discriminative and generative types, and propose two new confidence fusion strategies to improve calibration. Our experiments demonstrate that LLMs exhibit poorer calibration at the atomic level during long-form generation. More importantly, atomic calibration uncovers insightful patterns regarding the alignment of confidence methods and the changes of confidence throughout generation. This sheds light on future research directions for confidence estimation in long-form generation.
@@ 1 Introduction
While large language models (LLMs) (Touvron et al., 2023; Jiang et al., 2023; OpenAI, 2022) excel in various tasks, they still struggle with trustworthiness issues. LLMs often suffer from hallucinations, generating factually inaccurate content and misleading responses (Zhang et al., 2023b; Huang et al., 2023), which limits their application in high-risk real-world scenarios (Hu et al., 2023). To address this, confidence calibration aims to estimate the underlying uncertainty of model predictions and reflect the true likelihood of correctness (Guo et al., 2017). A calibrated model is crucial for real-world applications, as it allows us to determine the extent to which we can trust models’ predictions (Zhu et al., 2023; Mahaut et al., 2024). Improved calibration enables more reliable confidence estimation, warning users when not to trust the model and thus mitigating the impact of hallucinations.
@@
Most existing work on LLM calibration focuses on short-form QA tasks (Jiang et al., 2021; Tian et al., 2023; Zhu et al., 2023; Ulmer et al., 2024), using datasets like TriviaQA and Natural Questions (Joshi et al., 2017), where answers are typically under 10 words. In contrast, real-world queries often elicit much longer responses (Zhang et al., 2024; Yang et al., 2024b), spanning hundreds or thousands of words. In such cases, response quality is not simply binary, as answers may mix accurate and inaccurate statements.
Recent work has begun to address calibration in long-form generation (Zhang et al., 2024; Huang et al., 2024; Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024). Several approaches estimate a single confidence score for the entire response (macro calibration; upper, Figure 1), while others assess confidence at the level of atomic claims (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024) (atomic calibration; lower, Figure 1). However, previous work leaves several key research questions unanswered: Why is it important to evaluate calibration at the atomic-claim level? What factors influence calibration results at this level? What patterns can be observed by analyzing calibration at the atomic-claim level?
In this work, we systematically examine atomic calibration: A long response is decomposed into atomic claims, each containing a single factual statement, and confidence scores are assigned using various elicitation methods. To analyze different confidence elicitation methods, we categorize them into discriminative (intrinsic confidence estimation) and generative (external confidence assessment). Our experiments on three long-form QA datasets with seven LLMs reveal that: (1) Models that appear well-calibrated at the response level perform poorly at the atomic level (Figure 2, Table 1); (2) Leveraging atomic calibration enhances macro calibration (Table 2). These two reasons highlight the need for research on atomic-level calibration to develop better-calibrated models.
We further investigate the characteristics of discriminative and generative confidence. Our analysis yields two main findings: (1) Discriminative and generative methods are complementary; combining them improves calibration, while combinations within the same category offer limited gains. (2) Generative methods maintain consistent calibration across different model sizes, whereas discriminative methods benefit from increased model sizes. Motivated by finding (1), we propose two novel fusion strategies based on confidence agreement to integrate generative and discriminative confidence. Our strategies outperform existing fusion methods.
Our atomic-level analysis (Section 6) offers deeper insights into confidence method alignment and confidence changes during generation. Confidence methods within the same category align better, explaining why cross-category fusion is more effective. Interestingly, with discriminative methods, model confidence in atomic facts tends to decrease as generation progresses. In contrast, generative methods show the lowest average confidence in the middle of the generation process. These results highlight the necessity of fine-grained calibration evaluation for long-form generation, given us insights on model trustworthiness and usability.
2 Related Work
Atomic Claims Generation and Verification. Long-form responses often contain both correct and incorrect statements, which impact the overall factuality assessments. Min et al. (2023) propose breaking long responses into atomic facts and calculating the precision of these fact pieces to determine the overall factuality score. Wei et al. (2024) and Zhao et al. (2024) extend this paradigm by expanding the dataset to include more domains beyond biographies. Song et al. (2024) design VeriScore for diverse long-form generation tasks that feature both verifiable and unverifiable content. Chiang and Lee (2024) introduce D-FActScore, specifically designed for content with ambiguous entities. Decomposing long-form responses into atomic claims and fact-checking them individually has become a widely adopted pipeline.
Uncertainty and Calibration in Long-form Generations. Existing research on uncertainty estimation and calibration primarily focuses on multiple-choice or short-form questions (Zhu et al., 2023; Kuhn et al., 2022; Lin et al., 2023; Tian et al., 2023; Ulmer et al., 2024). There is an increasing interest on calibration for long-form generations. Huang et al. (2024) proposed a unified calibration framework for all text generation tasks, comparing distributions of both correctness and the associated confidence of responses. Band et al. (2024) introduced linguistic calibration, where models explicitly express their uncertainty during long-form generation. Zhang et al. (2024) proposed Luq, an uncertainty estimation method tailored to long-form generation, demonstrating its effectiveness in ensembling different LLMs. Another line of work (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024; Yuan et al., 2024) decomposes sentences into atomic claims and assigns confidence scores to each claim. However, a unified definition of atomic-level calibration remains lacking. Clarifying this concept and identifying key influencing factors are essential steps toward improving calibration in long-form generation.
@@ 3 Atomic Calibration
For a language model $\mathcal{M}$, let $x\sim M(x\mid q)$ denote the response generated by $\mathcal{M}$ for a query $q$, $x\in\mathcal{X}$. Let $y\in\mathcal{Y}_{t}$ be the corresponding label, representing a quality score ranging from 0 to 1 for a specific task $t\in T$. Unlike multiple-choice or short-form questions, which mainly assess correctness, tasks in $T$ cover diverse dimensions such as factuality, coherence, and creativity.
We define a probability prediction function $f:\mathcal{X}\to\Delta^{\lvert\mathcal{Y}_{t}\rvert}$, where $\Delta^{\lvert\mathcal{Y}_{t}\rvert}$ denotes the $\lvert\mathcal{Y}_{t}\rvert$-dimensional probability simplex. Here, $f(x)_{y}$ represents the probability assigned to label $y$ for a generated output $x$. In this work, we focus on calibrating factuality, as hallucinations are a well-known issue in LLMs (Zhang et al., 2023b; Huang et al., 2023), and the factuality of atomic claims can be assessed objectively. In this setting, $\mathcal{Y}$ denotes $\mathcal{Y}_{t}$ for the factuality task $t$, where $\mathcal{Y}\subseteq[0,1]$ reflects the factuality level of a response. Following Guo et al. (2017), we define the calibration of each response as follows:
@@
$\mathbb{P}(y\mid f(x)_{y}=\beta)=\beta,\quad\forall\beta\in[0,1].$
@@
$\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta_{i},\quad\forall\beta_{i}\in[0,1].$
@@ 4 Confidence Elicitation Methods
In this section, we define two types of confidence elicitation methods: generative and discriminative. We then introduce two novel confidence fusion strategies that considers confidence agreement when combining confidence scores. For the response $x$ to a query $q$, $x$ is broken into atomic claims $C$. Following previous work (Min et al., 2023; Wei et al., 2024; Zhao et al., 2024), each atomic claim contains a single piece of information and must be self-contained. For generative methods, we sample an additional set of responses $K$, and compare them against the original response $x$. For each atomic claim in $C$, we assign it a confidence score.
@@ 4.1 Generative Methods
Generative methods assume that the consistency between different generation samples provides a reliable estimation of model uncertainty (Zhang et al., 2024; Jiang et al., 2024). Generally, an additional natural language inference (NLI) model is used to calculate the consistency. In particular, we have the following two variations:
Gen-Binary. The basic assumption is that if a fact is frequently conveyed when sampled multiple times, the model is considered “confident” about that fact. For an atomic claim $c_{i}$ in $C$, we utilize a NLI model $\mathcal{M}_{\mathrm{NLI}}$ to examine whether $c_{i}$ is supported or not supported by each of the additional samples. Let $K_{s}$ be the set of samples supporting $c_{i}$. Then, the confidence in $c_{i}$ is calculated as
$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K\rvert}.$
Gen-Multi. Gen-Multi assumes that the model is more confident in facts that are consistently expressed. Unlike Gen-Binary, it further divides the “not supported” ($K_{ns}$) into “conflict” ($K_{c}$) if the fact is presented differently in the sample, and “not mentioned” ($K_{nm}$) if the fact is not mentioned in the sample. We then calculate the confidence by only considering supporting and conflicting samples:
$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K_{s}\rvert+\lvert K_{c}\rvert}.$
@@ 4.2 Discriminative Methods
Discriminative methods assess uncertainties by asking the model itself (Tian et al., 2023; Xiong et al., 2023). This is motivated by the findings that models tend to perform better on discriminative tasks (Saunders et al., 2022), and thus they may already possess the capability to estimate the confidence of their own outputs in a discriminative manner.
Dis-Single. Following Kadavath et al. (2022); Tian et al. (2023), we directly ask the model whether one single atomic claim is true or false. The probability the model assigns to token “True” ($P(true)$) in its generation is viewed as the confidence. As each atomic claim is judged individually, one advantage of this method is that there is no cross-claim influences when the model makes confidence judgments.
@@
Dis-Rating. Instead of using $P(true)$, in Dis-Rating, we directly prompt the model to assign a numerical value representing its confidence in the atomic claim $c_{i}$. A score of 0 indicates no confidence, while 10 represents maximum confidence. An alternative approach is to use semantic expressions ranging from “Very Uncertain” to “Very Confident”. However, Tian et al. (2023) demonstrate LLMs achieve comparable or even better results using numerical values.
@@ 4.3 Confidence Fusion Strategies
Combining confidence scores has proven effective for calibration (Huang et al., 2024; Rivera et al., 2024), but existing methods typically only use a single fixed weight, $\alpha$, to combine the scores, ignoring the confidence disagreement. For instance, Rivera et al. (2024) computes the weighted average (WAvg) for confidence scores $A$ and $B$: $C=A\cdot\alpha+B\cdot(1-\alpha)$. This approach does not account for the agreement between the two scores. For example, when $\alpha=0.5$, confidences of 0 and 1 are treated the same as confidences of 0.4 and 0.6, although the former may indicate higher uncertainty due to a larger disagreement. To address this, we propose two simple but effective methods that consider confidence disagreement $d=B-A$.
AdjustedAlpha adjusts the weight $\alpha$ based on the confidence difference:
$\alpha^{\prime}=\alpha+\gamma_{\text{a}}\cdot d,$
where $\gamma_{\text{a}}$ is a small constant (e.g., 0.1), followed by $C^{\prime}=A\cdot\alpha^{\prime}+B\cdot(1-\alpha^{\prime})$.
DampedFusion applies a damping factor based on the agreement:
$\gamma(d)=1-k\cdot|d|,$
where $k$ is a small constant (e.g., 0.02) that controls the damping sensitivity, followed by $C^{\prime}=C\cdot\gamma(d)$. For baselines, we also include: MinConf, which selects the minimum confidence; HMean, which calculates the harmonic mean; and ProdConf, which multiplies the confidences.
@@ 5.1 Experiment Setup
Models. We utilize seven LLMs from three model families with varying sizes: Llama3 Instruct (8B and 70B) (Meta, 2024), Mistral Instruct (7B and 8x7B) (Jiang et al., 2023), and Qwen2 Instruct (7B, 52B-A14B, and 72B) (Yang et al., 2024a).
@@
Atomic Facts Generation and Verification. For all three datasets, we apply a FActScore-based (Min et al., 2023) factuality assessment approach. We first use GPT-4o to decompose the entire response into atomic facts. These atomic facts are then verified using GPT-4o, cross-referenced with evidence from Wikipedia and Google Search. The detailed prompts for generating atomic facts are provided in Appendix I.
Confidence Elicitation. We use P(true) (Kadavath et al., 2022), Self-Rating (Tian et al., 2023), Semantic Entropy (SE) (Kuhn et al., 2022), and Sum of Eigenvalues (EigV) (Lin et al., 2023) as the baseline confidence elicitation methods. They are all calculated in response-level. For Gen-Binary, we apply the Llama-3-8B-Instruct for better NLI performance. For WAvg, AdjustedAlpha, and DampedConf, we use a separate validation set for hyper-parameter tuning.
Metrics. We use Expected Calibration Error (ECE) (Naeini et al., 2015) and Brier Score (BS) (Brier, 1950) as the primary metrics. These metrics are applicable to both atomic and macro calibration (see details in Appendix A), enabling a direct comparison between them. Additionally, we include AUROC to evaluate atomic calibration and Spearman Correlation for a more instance-specific assessment in macro calibration.
Figure 2: Comparison of atomic level and response-level calibration for ECE and Brier Score. Atomic-level performance is generally worse than response-level performance, with data points consistently lying above the identity line.
@@
ECE $\downarrow$ BS $\downarrow$ SC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ SC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ SC $\uparrow$
@@
P(true) 45.1 25.9 30.2 16.3 4.8 18.9 25.7 13.5 40.5
Self-Rating 38.7 23.4 40.5 14.1 4.2 21.5 18.6 12.9 50.2
SE 37.4 21.8 42.1 13.5 3.4 23.0 17.8 11.7 52.0
EigV 36.8 21.2 43.0 13.0 3.2 23.8 17.2 11.3 53.0
@@
Dis-Rating 25.7 11.7 73.8 2.9 1.6 34.1 3.6 3.5 71.7
Dis-Single 27.2 13.7 58.0 8.7 2.6 20.9 14.4 7.3 55.9
Gen-Binary 5.6 3.3 79.8 3.0 1.1 52.7 7.8 4.6 70.0
Gen-Multi 35.8 18.1 71.4 11.6 2.7 37.5 22.0 10.5 62.6
@@
P(true) 44.5 27.1 32.8 16.7 7.4 22.0 24.3 19.8 41.2
Self-Rating 37.1 26.4 42.3 14.5 6.5 26.1 18.1 14.5 52.0
SE 36.5 23.9 44.1 13.8 3.7 28.0 17.4 14.3 53.4
EigV 35.9 23.3 45.3 13.3 3.5 36.5 16.9 13.9 54.4
@@
Dis-Rating 41.4 22.7 55.0 4.4 1.7 40.8 16.9 9.8 60.4
Dis-Single 16.0 6.6 70.3 8.8 3.1 32.8 10.4 6.5 65.3
Gen-Binary 8.5 3.8 74.9 2.5 1.0 64.1 10.3 5.0 73.9
Gen-Multi 38.7 20.1 60.7 11.9 2.8 49.6 26.2 13.4 65.6
@@
P(true) 45.0 27.9 33.5 11.2 5.6 28.3 12.7 15.6 35.4
Self-Rating 24.3 25.1 48.2 6.9 4.9 36.7 9.8 14.7 48.0
SE 22.9 23.1 49.8 6.5 3.7 38.9 8.9 13.9 49.2
EigV 22.4 22.5 50.7 6.2 3.5 39.8 8.5 13.5 50.2
@@
Dis-Rating 40.7 21.3 63.0 4.4 1.9 29.9 9.1 6.3 54.0
Dis-Single 19.8 8.4 52.8 4.9 2.4 30.9 5.3 4.8 60.4
Gen-Binary 5.4 3.2 72.4 2.0 0.9 67.6 6.5 3.2 72.2
Gen-Multi 38.8 20.0 43.1 11.4 2.6 52.1 21.6 9.5 63.2
Table 2: Macro Calibration Results. All the numbers are in percentages.